Solve the system of equations by graphing.
3x-2y=0
-3x-y=-18

Answer by richwmiller(17219)

To solve by graphing you must graph it. Do you grasp that? (That is a pun.)
for each equation
let x=0 then find y
let y=0 then find x
Do them separately and draw the line then do the other.
Where the lines cross is the "solution"
This is not required for your homework and should not be handed in with it.
To find the "solution" algebraically add the two equations
-3y=-18
y=6
-3x-y=-18
-3x-6=-18
-3x=-12
x=4
(4,6)